# spade: Synthesizing Assertions for Large Language Model Pipelines

### v1

@@ 2.3. From Taxonomy to Assertions
Construct $\Delta\mathcal{P}_{\_}5$ “- Include elements from the movie’s genre, cast, and themes that align with the user’s interests. + Mention the movie’s genre and any shared cast members…” Prompt LLM with $\Delta\mathcal{P}_{\_}5$ to identify the delta type(s) and assertion criteria based on our taxonomy Synthesize Function(s)
Figure 3. Generating candidate assertion functions from a $\Delta\mathcal{P}$.
Using our taxonomy, spade employs an LLM to categorize and generate assertions for prompt deltas, detailed in Appendix C. For each $\Delta\mathcal{P}_{\_}i$, spade prompts the LLM to suggest as many criteria as possible for assertions, each aligning with a taxonomy category (e.g., Figure 3). A criterion is loosely defined as some natural language expression that operates on an example and evaluates to True or False (e.g., “check for conciseness”). spade analyzes every $\Delta\mathcal{P}_{\_}i$ instead of just last prompt version for several reasons: developers often remove instructions from prompts to reduce costs while expecting the same behavior (Parnin et al., 2023), prompts contain inherent ambiguities and imply multiple ways of evaluating some criteria, and complex prompts may lead to missed assertions if only one version is analyzed. Consequently, analyzing each $\Delta\mathcal{P}_{\_}i$ increases the likelihood of generating relevant assertions.
For each delta, spade collects the criteria identified and prompts the LLM again to create Python assertion functions. The synthesized functions can use external Python libraries or pose binary queries to an LLM for complex criteria. For function synthesis, the LLM is instructed that if the criterion is vaguely specified or open to interpretation, such as “check for conciseness,” it can generate multiple functions that each evaluate the criterion. In this conciseness example, the LLM could return multiple functions—a function that splits the response into sentences and ensures that there are no more than, say, 3 sentences, a function that splits the response into words and ensures that there are no more than, say, 25 words, or a function that sends the response to an LLM and asks whether the response is concise. The overall outcome of this step is a multiset of candidate functions $F=\{f_{\_}1,\dots,f_{\_}m\}$.
@@
We adopt a two-step process in our approach, as it has been demonstrated that breaking tasks into steps can enhance LLM accuracy (Wei et al., 2022; Wu et al., 2022b; Grunde-McLaughlin et al., 2023). However, it’s worth noting that with daily advancements in LLMs, a one-step process might already be feasible. Additionally, while our taxonomy guides LLM-generated assertions now, future LLMs may implicitly learn these categories through reinforcement learning from human feedback (Christiano et al., 2017). Nevertheless, knowing the taxonomy-based categories associated with candidate assertions may help in filtering them.
2.4. LangChain Deployment
@@
To assess the potential of our auto-generated assertions, in the first week of November 2023, we released an early prototype of spade’s candidate assertions via a Streamlit application [^4] https://spade-beta.streamlit.app/. At the time of public release, the Streamlit application had a different set of prompts to generate assertions; these prompts generated relatively few assertions compared to the number of assertions generated in this paper’s version of spade. However, the taxonomy has remained the same. In the Streamlit app, a developer can either paste their prompt template that they want to generate assertions for, or they can point to their LangChain Hub prompt template (which automatically contains prompt version history via commits). The app then visualizes the identified taxonomy categories in the user’s prompt and displays the candidate assertions, as shown in the screenshot in Figure 4.
@@
We found significant interest in auto-generated LLM pipeline assertions: there were over 1300 runs of the app for custom prompt templates (i.e., not the sample default prompt template in app) within two weeks. These runs span a variety of fields, including medicine, education, cooking, and finance (Figure 5). Users for 10 of the 1300+ runs gave an optional “thumbs up” feedback on the assertions generated for them, and users for 43 of these runs clicked the “download assertions” button, which downloads the candidate assertions as a Python file. We note that users can copy the code in the assertions instead of downloading the assertions as a Python file, and we were unable to measure the copy events. No users clicked the “thumbs down” button—which is not to say that the generated assertions were perfect. In fact, anecdotally, we have seen the candidate assertion quality improve over the last several months as GPT-4 continually improves. On average, 3.3 assertions were generated per run, with minimum 1 and maximum 10 assertions generated for a run. Most of the runs corresponded to a single prompt version. We also found a deployment of spade-generated assertions from a LangChain user who built a “chat-with-your-pdf” tool [^5] https://twitter.com/thcalafatidis/status/1728144652119769394: in their words, “When I saw it I didnt beleive it could work that well, but it really did and made the evaluation process fun and ez.”
@@
[…]
@@
In LLM pipelines with numerous prompt versions, we observed two main patterns. First, prompt engineering often leads to many similar assertions, and redundancy can be a headache at deployment when a developer has to keep track of so many assertions. For instance, a pipeline to summarize lecture transcripts [^6] https://smith.langchain.com/hub/kirby/simple-lecture-summary had 14 prompt versions, with many edits localized to the paragraph providing instructions on titles, speakers, and dates, creating multiple overlapping assertions. 10 assertions assessed the summary’s thesis; two are as follows:
Second, many assertions may be incorrect, causing runtime errors or false failures. Assessing the accuracy of these assertions is challenging, particularly for complex or LLM-invoking ones, where even experienced developers might not be able to gauge their effectiveness without viewing the results of the functions on many examples. Since there can be 50+ assertions generated for just a handful of prompt versions (as demonstrated in Section 4), manually filtering them by eyeballing failure rates for each subset of assertions is impractical. Therefore, we adopt an automated approach for filtering.
@@ 3. Filtering Candidate Assertions
Here, we address the issue of redundant and incorrect assertions identified in Section 2.4, particularly in pipelines with numerous prompt versions. Filtering this candidate set not only improves efficiency when deploying the assertions to run in production, but it also reduces cognitive overhead for the developer.
@@ 3.1. Definitions
Consider $e_{\_}i$ as an example end-to-end execution (i.e., run) of an LLM pipeline on some input. Let $E$ be the set of all such example runs (this set is not provided upfront, as we will deal with shortly). We define an assertion function $f:E\rightarrow\{0,1\}$, where 11 11 indicates success and 00 0 indicates failure. Let $F^{\prime}=\{f_{\_}1,f_{\_}2,\ldots,f_{\_}k\}$ be a set of $k$ assertions. An example $e_{\_}i$ is deemed successful by this set if and only if it satisfies all assertions in $F^{\prime}$. Specifically:
$\hat{y}_{\_}i=\begin{cases}1&\text{if }f(e_{\_}i)=1\quad\forall f\in F^{\prime},\\ 0&\text{otherwise.}\end{cases}$
We denote $y_{\_}i\in\{0,1\}$ to represent whether the LLM chain developer considers $e_{\_}i$ to be a success (1) or failure (0). Given all $m$ candidate assertions $F={f_{\_}1,f_{\_}2,\ldots,f_{\_}m}$, the objective is to select $F^{\prime}\subseteq F$ such that $\hat{y}_{\_}i=y_{\_}i$ for most examples in $E$, with $F^{\prime}$ being as small as possible. This goal involves maximizing failure coverage and minimizing the false failure rate and selected function count, expressed as follows:
Definition 3.1 ().
@@
$\text{Coverage}\left(F^{\prime}\right)=\frac{\sum_{\_}{i}\mathbb{I}\left[\hat{y}_{\_}i=0\land y_{\_}i=0\right]}{\sum_{\_}{i}\mathbb{I}\left[y_{\_}i=0\right]}$
Definition 3.2 ().
@@
$\text{FFR}\left(F^{\prime}\right)=\frac{\sum_{\_}{i}\mathbb{I}\left[\hat{y}_{\_}i=0\land y_{\_}i=1\right]}{\sum_{\_}{i}\mathbb{I}\left[y_{\_}i=1\right]}$
In both Definition 3.1 and Definition 3.2, $\hat{y}_{\_}i$ represents set $F^{\prime}$’s prediction for the $i$-th example, $y_{\_}i$ is the actual outcome determined by the LLM chain developer, while $\mathbb{I}$ represents the indicator function. In practice,  Definition 3.1 and Definition 3.2 are impossible to compute since the universe of examples $E$ is unknown. So, for now, we assume access only to a subset $E^{\prime}\subset E$ of labeled LLM responses, where $E^{\prime}$ is a manually provided set of example runs and may not contain all the types of failures the LLM pipeline could hypothetically observe—an issue we will deal with in Section 3.3. Thus, we replace Definition 3.1 and Definition 3.2 with $\text{Coverage}_{\_}{E^{\prime}}\left(F^{\prime}\right)$ and $\text{FFR}_{\_}{E^{\prime}}\left(F^{\prime}\right)$, omitting the subscript $E^{\prime}$ for brevity. Table 7 in Appendix A summarizes the notation used throughout this section.
@@ 3.2. Coverage Problem Formulation
The above problem may be infeasible with certain values of $\alpha$, for multiple reasons. To see this, consider the case where no assertion catches a failure while $\alpha=1$. We ignore this (and similar cases) for now and defer their discussion to Section 4. To expand out the above, we introduce a matrix $M$ (size $n\times m$) to track each assertion’s result on each example $e_{\_}i$, where $M_{\_}{ij}=1$ if $f_{\_}j(e_{\_}i)=1$ and $M_{\_}{ij}=0$ otherwise. We also define binary variables $x_{\_}j$ and $w_{\_}{ij}$ to represent whether an assertion is chosen and if it marks an example as a failure: $w_{\_}{ij}=\left(1-M_{\_}{ij}\right)\cdot x_{\_}j$, which is based on whether $f_{\_}j$ denotes $e_{\_}i$ as a failure and $f_{\_}j$ is included in $F^{\prime}$. We additionally introduce the binary variable $u_{\_}i$ to represent whether a failed example is covered by any selected assertion:
$\displaystyle u_{\_}i\leq\sum_{\_}{j=1}^{m}w_{\_}{ij},\quad\forall i\in\left[1,n\right]:y_{\_}i=0.$
@@
The product $w_{\_}{ij}=(1-M_{\_}{ij})\cdot x_{\_}j$ in the numerator is 1 if $f_{\_}j$ is included in $F^{\prime}$ and incorrectly marks a good example as a failure. The $\max$ function ensures that as long as there is some selected function $f_{\_}j$ that marks $e_{\_}i$ as a failure, the product is 1. Multiplying $y_{\_}i$ by the $\max$ result ensures that the numerator is incremented only when $y_{\_}i=1$, or when the example $e_{\_}i$ is actually not a failure. This numerator is thus equivalent to the numerator in Definition 3.2, i.e., $y_{\_}i=1\land\hat{y}_{\_}i=0$. The entire fraction is less than the threshold $\tau$, which represents the maximum allowed false failure rate across the examples. To formulate FFR according to (1), we introduce a new binary variable $z_{\_}i$, which defines whether $F^{\prime}$ denotes $e_{\_}i$ as a failure while $e_{\_}i$ is actually a successful example (i.e., false failure):
$\displaystyle z_{\_}i\geq y_{\_}i\cdot w_{\_}{ij},\quad\forall i\in\left[1,n\right];\,\forall j\in\left[1,m\right].$
@@
$\displaystyle\quad z_{\_}i\geq y_{\_}i\cdot w_{\_}{ij},\quad\forall i\in\left[1,n\right],\ \forall j\in\left[1,m\right];\quad\frac{\sum_{\_}{i=1}^{n}z_{\_}i}{\sum_{\_}{i=1}^{n}\mathbb{I}\left[y_{\_}i=1\right]}\leq\tau.$
We refer to a solution for this ILP as $\textsc{spade}_{\_}{\text{cov}}$. Trivially, the problem is NP-hard for $\tau=0$ and $\alpha=1$, via a simple reduction from set cover, and is in NP, since it can be stated in ILP form.

### v2

@@ 2.3. From Taxonomy to Assertions
$\spadesuit$ spade Construct $\Delta\mathcal{P}_{\_}5$ “- Include elements from the movie’s genre, cast, and themes that align with the user’s interests. + Mention the movie’s genre and any shared cast members…” Prompt LLM with $\Delta\mathcal{P}_{\_}5$ to identify the delta type(s) and assertion criteria based on our taxonomy Synthesize Function(s)
Figure 3. Generating candidate assertions from a $\Delta\mathcal{P}$.
@@
As we saw previously, prompt delta as categorized into nodes in our taxonomy often correspond to meaningful assertion criteria. Next, we need a method to automatically synthesize the data quality assertions from the prompt deltas. A natural idea is to prompt an LLM to generate assertion functions corresponding to relevant categories in our taxonomy given the prompt deltas. We tried this approach with GPT-4 in January 2024 and observed several omitted assertions that clearly corresponded to categories in our taxonomy. Therefore, we adopt a two-step prompting process in our approach, as it has been demonstrated that breaking tasks into steps can enhance LLM accuracy (Wei et al., 2022; Wu et al., 2022b; Grunde-McLaughlin et al., 2023). In the first step, we prompt GPT-4 for natural language descriptions of assertion criteria, and in the second step, we prompt GPT-4 to generate Python functions that implement such criteria. However, it’s worth noting that with daily advancements in LLMs, a one-step process might already be feasible. Additionally, while our taxonomy guides LLM-generated assertions now, future LLMs may implicitly learn these categories through reinforcement learning from human feedback (Christiano et al., 2017). Nevertheless, knowing the taxonomy-based categories associated with candidate assertions may help in filtering them.
More specifically, for each prompt delta $\Delta\mathcal{P}_{\_}i$, we first prompt an LLM to suggest as many criteria as possible for assertions—each aligning with a taxonomy category from Figure 3. A criterion is loosely defined as some natural language expression that operates on a given output or example and evaluates to True or False (e.g., “check for conciseness”). Our method analyzes every $\Delta\mathcal{P}_{\_}i$ instead of just the last prompt version for several reasons: developers often remove instructions from prompts to reduce costs while expecting the same behavior (Parnin et al., 2023), prompts contain inherent ambiguities and imply multiple ways of evaluating some criteria, and complex prompts may lead to missed assertions if only one version is analyzed. Consequently, analyzing each $\Delta\mathcal{P}_{\_}i$ increases the likelihood of generating relevant assertions.
For each delta, our method collects the criteria identified and prompts an LLM again to create Python assertion functions. The synthesized functions can use external Python libraries or pose binary queries to an LLM for complex criteria. For function synthesis, the LLM is instructed that if the criterion is vaguely specified or open to interpretation, such as “check for conciseness,” it can generate multiple functions that each evaluate the criterion. In this conciseness example, the LLM could return multiple functions—a function that splits the response into sentences and ensures that there are no more than, say, 3 sentences, a function that splits the response into words and ensures that there are no more than, say, 25 words, or a function that sends the response to an LLM and asks whether the response is concise. The overall outcome of this is a multiset of candidate functions $F=\{f_{\_}1,\dots,f_{\_}m\}$. The two prompts for generating assertion criteria and functions can be found in Appendix C.
2.4. Initial Deployment
@@
To assess the potential of our auto-generated assertions, in November 2023, we released an early prototype of spade’s candidate assertion generation framework via a Streamlit application [^3] https://spade-beta.streamlit.app/. At the time of public release, the Streamlit application had a different set of prompts to generate assertions; these prompts generated relatively few assertions compared to the number of assertions generated in this paper’s version of spade. However, the taxonomy has remained the same. In the Streamlit app, a developer can either paste their prompt template that they want to generate assertions for, or they can point to their LangChain Hub prompt template (which contains prompt version histories via commits). The app then visualizes the identified taxonomy categories in the user’s prompt and displays the candidate assertions, as shown in the screenshot in Figure 4.
@@
Tool Usage Insights and Feedback. From the reception to our Streamlit application, we found significant interest in auto-generated data quality assertions for LLM pipelines: there have been over 2000 runs of the app for custom prompt templates (i.e., not the sample default prompt template in app). These runs span many fields, including medicine, education, cooking, and finance—providing insights from a diversity of use cases for LLM pipelines. Figure 5 shows a rough breakdown of the use cases people wanted to generate assertions for; however, it’s important to note that some runs may not cleanly fit into a single category (e.g., a chatbot for telehealth-related questions for a medical provider could be in “customer support” and “health”). Interestingly, 8% of tasks related to conversational assistants, and we observed instances of at least four different companies generating assertions for their chatbots, given that the company name was in the prompt. Users for 48 of these runs clicked the “download assertions” button, which downloads the candidate assertions as a Python file. We note that users can also directly copy the assertions code displayed, instead of downloading the assertions as a Python file, and we did not measure the copy events. No users clicked the “thumbs down” button—which is not to say that the generated assertions were perfect. In fact, anecdotally, we have seen the candidate assertion quality improve over the last several months as GPT-4 continually improves. On average, 3.3 assertions were generated per run, with minimum 1 and maximum 10 assertions generated for a run. Most of the runs corresponded to a single prompt version. After our Streamlit app was released, we found an unprompted review of spade-generated assertions from a LangChain user who built a “chat-with-your-pdf” tool [^4] https://twitter.com/thcalafatidis/status/1728144652119769394: in their words, “When I saw it I didnt beleive it could work that well, but it really did and made the evaluation process fun and ez.”
Observations about Assertion Criteria. Across the 2000+ runs, with regards to our taxonomy, assertion criteria were most commonly derived from inclusion and exclusion instructions. For example, for a shopping assistant, where the objective was to find the most relevant product related to a customer’s query, responses were required to include “all the features [the customer] asked for.” In another customer support agent example, responses were required to include “exact quotes in the [context] relevant to the [customer’s] question…word for word.” One common exclusion instruction across chat-related tasks was to avoid any discussion unrelated to the end-user’s question; however, we noticed that GPT-4 struggled to generate assertion criteria around such a generic exclusion instruction. We found Python functions generated to implement criteria such as “avoid unrelated ideas”, did not use an LLM to check the criteria. Instead, they checked for the presence of specific phrases like “unrelated” and “I’m sorry, I did not find anything related” in the response. Such errors indicate the limitations of current state-of-the-art LLMs and may suggest the need for specialized, fine-tuned LLMs for generating assertions in future work; here, we work with the limitations of present-day state-of-the-art LLMs. We discuss this further in Section 4.4.
A new pattern we noticed after deployment, which wasn’t seen in our initial analysis, is that developers often wish to hide certain parts of the LLM workflow in the outputs. This could be viewed as a special instance of the exclusion category in our taxonomy. For instance, in a prompt related to enterprise automation, the first instruction was to write a query targeting a specific database (e.g., “Write a SQL query to fetch the most relevant table from the MySQL database”). However, another instruction in the prompt specified that the name of the database should not appear in the final summary returned to the end-user (e.g., “Do not mention that you queried the MySQL table X”). We were pleasantly surprised to find that our process for generating criteria successfully identified and implemented such exclusion instructions accurately.
@@
$\spadesuit$ spade 00 055 551010 10101515 15152020 2020 Health IT and Programming Project Management Customer Support Education Sports and Entertainment Cooking Finance Marketing Others 20.820.8 20.820.816.716.7 16.716.712.512.5 $12.5$ 12.512.58.38.3 $8.3$ $8.3$ 8.38.34.24.2 $4.2$ $4.2$ 4.24.2Percentage
@@
Observations about Assertions. In LLM pipelines with numerous prompt versions, we observed two main patterns. First, prompt engineering often leads to many similar assertions, and redundancy can be a headache at deployment when a developer has to keep track of so many assertions. For instance, a pipeline to summarize lecture transcripts [^5] https://smith.langchain.com/hub/kirby/simple-lecture-summary had 14 prompt versions, with many edits localized to the paragraph providing instructions on titles, speakers, and dates, creating multiple overlapping assertions. 10 assertions assessed the lecture’s central thesis; two are as follows:
$\spadesuit$ spade Second, many assertions may be incorrect, causing runtime errors or false failures. Assessing the accuracy of these assertions is challenging, particularly for those that are complex or invoke LLMs themselves, where even experienced developers might not be able to gauge their effectiveness without viewing the results of the functions on many examples. Even if the assertions are not incorrect, they still may have undesirable failure rates or coverage, which is often hard for a developer to reason about in conjunction with other assertions. Since there can be 50+ assertions generated for just a handful of prompt versions (as demonstrated in Section 4), manually filtering them by eyeballing failure rates for each subset of assertions is impractical. Therefore, we adopt an automated approach for filtering, as discussed in the following section.
@@ 3. Filtering Candidate Assertions
As we saw in the previous section, we often have redundant and incorrect assertions, particularly in pipelines with numerous prompt versions. Here, we focus on filtering this candidate set to a smaller number, which not only improves efficiency and reduces cost when deploying the assertions to run in production, but it also reduces cognitive overhead for the developer.
@@ 3.1. Definitions
Consider $e_{\_}i$ as an example end-to-end execution (i.e., run) of an LLM pipeline on some input. For the purposes of this section, we assume the prompt template $\mathcal{P}$ is fixed to be the final prompt version that the developer eventually decided on. We denote $y_{\_}i\in\{0,1\}$ to represent whether the developer considers $e_{\_}i$ to be a success (1) or failure (0).
Let $E$ be the set of all such example runs (this set is not provided upfront, as we will deal with shortly). We define an assertion function $f:E\rightarrow\{0,1\}$, where 11 11 indicates that the example was processed successfully by the LLM and 00 0 otherwise. Let $F^{\prime}=\{f_{\_}1,f_{\_}2,\ldots,f_{\_}k\}$ be a set of $k$ data quality assertions. An example $e_{\_}i$ is deemed successful by $F^{\prime}$ if and only if it satisfies all assertions in $F^{\prime}$. Specifically:
$\hat{y}_{\_}i=\begin{cases}1~{}~{}\text{if }f(e_{\_}i)=1\quad\forall f\in F^{\prime},&\text{ ~{}~{}~{}~{} // Deemed successful by all of }F^{\prime}\\ 0~{}~{}\text{otherwise.}&\text{ ~{}~{}~{}~{} // Deemed a failure by}\geq 1\text{ of }F^{\prime}\end{cases}$
Given all $m$ candidate assertions $F=\{f_{\_}1,f_{\_}2,\ldots,f_{\_}m\}$, the objective is to select $F^{\prime}\subseteq F$ such that $\hat{y}_{\_}i=y_{\_}i$ for most examples in $E$, with $F^{\prime}$ being as small as possible. This goal involves maximizing failure coverage and minimizing the false failure rate and selected function count, expressed as follows:
Definition 3.1.
@@
$\text{Coverage}\left(F^{\prime}\right)=\frac{\sum_{\_}{i}\mathbb{I}\left[\hat{y}_{\_}i=0\land y_{\_}i=0\right]}{\sum_{\_}{i}\mathbb{I}\left[y_{\_}i=0\right]}\text{ ~{}~{}~{}~{} // Failures caught by }F^{\prime}$
Definition 3.2.
@@
$\text{FFR}\left(F^{\prime}\right)=\frac{\sum_{\_}{i}\mathbb{I}\left[\hat{y}_{\_}i=0\land y_{\_}i=1\right]}{\sum_{\_}{i}\mathbb{I}\left[y_{\_}i=1\right]}\text{ ~{}~{}~{}~{} // Non-failures flagged by }F^{\prime}$
In both definitions above, $\hat{y}_{\_}i$ represents set $F^{\prime}$’s prediction for the $i$-th example, $y_{\_}i$ is the actual outcome, while $\mathbb{I}$ is the indicator function. In practice, coverage and FFR are impossible to compute since the universe of examples $E$ is unknown. So, for now, we assume access to a subset $E^{\prime}\subset E$ of labeled LLM responses, where $E^{\prime}$ is a manually provided set of example runs and may not contain all the types of failures the LLM pipeline could observe—an issue we will deal with in Section 3.3. Thus, we replace Definition 3.1 and Definition 3.2 with $\text{Coverage}_{\_}{E^{\prime}}\left(F^{\prime}\right)$ and $\text{FFR}_{\_}{E^{\prime}}\left(F^{\prime}\right)$, omitting the subscript $E^{\prime}$ for brevity. Table 8 in Appendix A summarizes the notation used throughout this section.
@@ 3.2. Coverage Problem Formulation
The above problem may be infeasible with certain values of $\alpha$, for multiple reasons. To see this, consider the case where no assertion catches a specific failure, while $\alpha$ is set to 11 11. We ignore this (and similar cases) for now and defer their discussion to Section 4. To expand out the above, we introduce a matrix $M$ (size $n\times m$) to track each assertion’s result on each example $e_{\_}i$, where $M_{\_}{ij}=1$ if $f_{\_}j(e_{\_}i)=1$ (i.e., $F_{\_}j$ deems $e_{\_}i$ a success) and $M_{\_}{ij}=0$ otherwise. We also define binary variables $x_{\_}j$ and $w_{\_}{ij}$ to represent whether an assertion is chosen and if it marks an example as a failure: $w_{\_}{ij}=\left(1-M_{\_}{ij}\right)\cdot x_{\_}j$, which is based on whether $f_{\_}j$ denotes $e_{\_}i$ as a failure and $f_{\_}j$ is included in $F^{\prime}$. We additionally introduce the binary variable $u_{\_}i$ to represent whether a failed example is covered by any selected assertion:
$\displaystyle u_{\_}i\leq\sum_{\_}{j=1}^{m}w_{\_}{ij},\quad\forall i\in\left[1,n\right]:y_{\_}i=0.\text{~{}~{}~{}~{}~{}~{}~{}~{}// Coverage of failure }e_{\_}i$
@@
The product $w_{\_}{ij}=\left(1-M_{\_}{ij}\right)\cdot x_{\_}j$ in the numerator is 1 if $f_{\_}j$ is included in $F^{\prime}$ and incorrectly marks a good example as a failure. The $\max$ function ensures that as long as there is some selected function $f_{\_}j$ that marks $e_{\_}i$ as a failure, the product is 1. Multiplying $y_{\_}i$ by the $\max$ result ensures that the numerator is incremented only when $y_{\_}i=1$, or when the example $e_{\_}i$ is actually not a failure. This numerator is thus equivalent to the numerator in Definition 3.2, i.e., $y_{\_}i=1\land\hat{y}_{\_}i=0$. The entire fraction is less than the threshold $\tau$, which represents the maximum allowed false failure rate across the examples. To formulate FFR according to (1), we introduce a new binary variable $z_{\_}i$, which defines whether $F^{\prime}$ denotes $e_{\_}i$ as a failure while $e_{\_}i$ is actually a successful example (i.e., false failure):
$\displaystyle z_{\_}i\geq y_{\_}i\cdot w_{\_}{ij},\quad\forall i\in\left[1,n\right];\,\forall j\in\left[1,m\right].\text{~{}~{}~{}~{}~{}~{}~{}~{}// }e_{\_}i\text{ is a false failure}$
@@
$\displaystyle\quad x_{\_}j,u_{\_}i,z_{\_}i,w_{\_}{ij}\in\{0,1\},\quad\forall i\in\left[1,n\right],\ \forall j\in\left[1,m\right].$
We refer to a solution for this ILP as $\textsc{spade}_{\_}{\text{cov}}$. Trivially, the problem is NP-hard for $\tau=0$ and $\alpha=1$, via a simple reduction from set cover, and is in NP, since it can be stated in ILP form. In our case, given tens of candidate assertions and fewer than 100 examples $e_{\_}i$, the ILPs tend to be of reasonable size (i.e., thousands of variables). Most ILP solvers can efficiently and quickly such programs.